# Past, Present, and Future of Citation Practices in HCI

### v1

@@ 4. Findings
Over the course of 34 years (from 1981 to 2024), authors at the CHI conference cited a total of 558,080 articles. The mean number of references per CHI article over the entirety of the CHI proceedings is 48.35 references ($SD=35.09$, $Min=0$, $Max=496$). However, the field of HCI has evolved from its humble beginnings, and a more in-depth look at these statistics is warranted. In particular, we focus on the rising number of mean references per article in the following section.
4.1. The Rising Number of References in CHI Articles
We first provide an historical overview of citation trends in the CHI proceedings, from the conference’s beginning to the current trend. We then extrapolate this trend into the future to provide an outlook for the year 2030. This year is not too distant (only six years from the time of writing), and we believe focusing on the year 2030 in our prediction of future trends will – if the current trend is not broken – provide an accurate and realistic estimation of the future of citations at CHI.
@@ 4.1.1. Historical development
In the early years of the CHI conference, it was not uncommon for authors to submit an article without a single citation to other works. The trend of zero-reference articles continued with only few exceptions until the year 1997 (see Figure 1 and Figure 3). From 1998 to 2017, the minimum number of references per article has been positive and in the one-digit range. Since 2017, the minimum number of references has increased into the double-digits.
@@
[…]
Number of references per CHI article from 1981 to 2024.
Between 1981 and 2015, the mean number of references per CHI article showed only modest linear growth (cf. Figure 1). After 2015, a clear and sudden quantitative change in the CHI community’s citation practices took place (see Figure 1 and Figure 3). Clearly, the CHI community entered a novel trajectory at this point in time, characterized by a year-by-year increase in the mean number of references per article. While there were on average 28.8 references per article in the 2015 proceedings, the number of references per article more than tripled to an average of 88.6 references per article in 2024.
For about two decades since its inception, the maximum number of references per CHI article hovered at around 40. After 2015, the maximum number of references skyrocketed to up to 496 references in one article in 2022. The box plots in Figure 3 demonstrate that the maximum number of references per article has become more volatile in recent years. During the years of the COVID-19 pandemic (2020 and 2021), one can observe a slump in the maximum number and upper quartile of the number of references per CHI article. However, there is overall a clear tendency for CHI authors to include more references in their articles than before the year 2015, and this number is monotonously growing each year. The number of references per article between 2006 and 2015 can be approximated with linear regression to a function $y_{t+1}={m_{1}}y_{t}+b_{1}$, with $m_{1}=0.79$, $b_{1}=-1555.92$, and $y_{t}\in[2006,2015]$, with mean absolute error $MAE=0.50$ and mean squared error $MSE=0.38$. The trend after 2015 is approximated by the function $y_{t+1}={m_{2}}y_{t}+b_{2}$, with $m_{2}=6.68$, $c=-13435.27$, and $y_{t}>2015$ (with $MAE=1.88$ and $MSE=6.16$).
To examine the differences between the linear regression models fitted to the two distinct periods, we employed a Chow Test, a test to determine whether there is a significant difference in the coefficients between two the linear regressions. The Chow Test indicates a substantial divergence in the regression parameters across the two periods under consideration ($F=145.47$, $p=4.30×e^{-10}$). This extremely low p-value strongly refutes the null hypothesis, suggesting that the regression parameters (slope and intercept) for the two periods are significantly distinct. This statistical evidence underscores the presence of a notable change in the trend between the two periods, corroborated by a clear shift in the slope. We assert that the observed variations in the dataset characteristics are not merely coincidental but indicative of a fundamental shift in the underlying dynamics across the two periods examined.
4.1.2. Extrapolation of the observed trend
We extrapolate the observed trend with the linear regression model and predict the mean number of references per article in future instances of the CHI conference proceedings (see Figure 4). We estimate that if the current observed trend is not broken, an article published at CHI 2030 will contain on average 129.5 references. The minimum number of references per article in 2030 is predicted to reach 33.7 references, which is higher than the mean number of references from the year 2015 (28.9 references).
Figure 4. Extrapolation of the trend in the mean and minimum number of references per CHI article. CHI articles are predicted to include, on average, almost 130 references in the year 2030. The minimum number of references per CHI article is estimated to reach 33.7 references per article, which is more than the 2015 average level of references.
[…]
Prediction of mean and minimum number of references per CHI article from 2024 to 2030.
4.2. Root Cause: A Change in Editorial Policies at the CHI Conference
A diverse and complex community consisting of thousands of members, such as the CHI community, is not likely to change its course suddenly, unless change is mandated from top down. In this section, we discuss the editorial change that can be considered the root cause for the observed trend.
@@ 4.2. Root Cause: A Change in Editorial Policies at the CHI Conference
Figure 5. Side-by-side comparison of excerpts from the CHI conference Call for Papers (CfPs) in 2015 and 2016 (own highlighting).
@@
Side-by-side comparison of excerpts from the CHI conference Call for Papers (CfPs) in the years 2015 and 2016.
4.2.1. Lifting the page restrictions in 2016 launched the CHI community on its expansive path
The CHI conference organizing committee (Program Chairs) in consultation with the Executive Committee of the ACM SIGCHI (Special Interest Group on Computer-Human Interaction) introduced a decisive change in the year 2016. In this year, the conference’s Call for Papers (CfP) moved away from counting references towards the page count limit, as highlighted in the excerpts reproduced in Figure 5. Articles published at CHI ’16 were allowed to include an unlimited number of references, as long as the article’s content fit the 10-page limit, with references no longer counting towards this page limit. While sub-committees were introduced as another significant change in the year 2016, allowing an unlimited number of references clearly is one of the main contributors to the increase in the mean number of references identified in our work. Further, the 2016 submission guide for CHI authors stressed the importance of citing relevant previous work, explicitly asking “is prior work adequately reviewed?” The 2016 guide highlighted the importance of reproducibility of authors work, with “letting others build on your work” being regarded as “the entire purpose of a CHI Paper.” These direct quotes from the 2016 guide were stark reminders and may have contributed to making the community pay special attention to their references in the year 2016. These policy changes nudged the community towards its expansive path. Awards presented to articles may have also played a role, as demonstrated in the following section.
4.2.2. Are high-citation articles rewarded in the CHI community?
Each year, awards are given to a select number of accepted articles at the CHI conference. Typically, between 14.6% and 24.6% of articles are given an award at the CHI conference. [^2] Based on awards presented from 2015–2024. These awards are a signal for the community on which articles are to be considered high quality. As a quality signal, awards can motivate authors to change their behavior. If articles with a high number of references were to systematically be presented with awards, this could potentially incentivize CHI authors to include more references in their articles. In this section, we examine whether there is such a systematic bias in awarded papers. We analyze the difference in the mean number of references in articles that received an award at the CHI conference compared to the mean number of references in articles that did not receive an award. As award, we consider both “best paper” awards and “honorable mentions.”
Table 1. Comparison of mean number of references in articles awarded a best paper award or honorable mention (Mawarded) with the mean number of references in unawarded articles (Mregular) in the CHI Proceedings from 2015 to 2024.
Mawarded Mregular t p
2015 31.2185 28.1534 2.2984 0.0220 *
2016 43.4696 40.7047 1.6156 0.1068
2017 49.8843 44.4969 2.7794 0.0056 **
2018 57.8651 49.8795 3.6289 0.0003 ***
2019 61.1419 54.9910 2.6155 0.0091 **
2020 66.0513 60.1246 2.6780 0.0076 **
2021 72.9648 70.9868 0.6910 0.4898
2022 87.9219 81.9332 1.5201 0.1290
2023 88.9688 84.7736 1.2218 0.2221
2024 92.9579 87.6505 1.7950 0.0729
@@ 4.2.2. Are high-citation articles rewarded in the CHI community?
Table 1 shows no evidence of a systematic bias towards awarding high-reference articles in recent proceedings. For instance, in the 2024 CHI Proceedings, the mean number of references for awarded articles was 92.96 versus 87.65 for articles without award. This means a high number of references is not considered a signal of high quality in the CHI community in recent years. However, between 2015 and 2020 (with exception of 2016), awarded articles had significantly more references than unawarded articles. The systematic award of articles with above-average references may have sent a quality signal to authors, thus potentially contributing to launching the CHI community on its current trajectory.
4.3. Additional Contextual Factors influencing HCI Citation Practices
The CHI community is vast and highly diverse, and there are many more factors that potentially shape and make up the citation practices of the members of the CHI community. While the editorial change introduced in 2016 is the primary reason for the change in the citation practices of HCI scholars, in this section, we take a step back and shed light on a number of broader contextual factors affecting the citation practices at the CHI conference. We do not claim that this list of factors is exhaustive. Beside the investigated changes, other latent factors could contribute to the observed trend, as is the case with any complex system under change. The investigated factors are an increase in the average number of authors (Section 4.3.1), an increase in the number of literature reviews (Section 4.3.2), an increasing number of citations to arXiv pre-print articles (Section 4.3.3), and an increase in citations to data and code repositories (Section 4.3.4).
@@ 4.3.1. Collaborative research is increasing
There has been an increase in collaborative research efforts in the field of HCI, leading to articles with a larger number of co-authors. Figure 6 depicts the number of authors per CHI article for each CHI proceeding. The mean number of authors per article has increased over time, from an average of 1.85 authors per CHI article in 1981 to an average of 5.17 authors per article in 2024.
Figure 6. Violin plot illustrating the distribution and central tendencies (mean and median) of the number of authors per CHI article from 1981 to 2024, highlighting the trend toward an increase in collaboration over time.
[…]
Number of authors per CHI article from 1981 to 2024.
Figure 7. Mean number of references per author in CHI articles from 1981 to 2024.
[…]
Mean number of references per author in CHI articles from 1981 to 2024.
@@
Figure 7 plots the mean number of references per author in CHI articles since the inception of the CHI Conference. In 1981, a CHI article would include on average 5.6 references per author. In 2024, this number has increased to 21.4 references per author. We note that while the mean number of references per author has seen a slow increase over time, there is a significant change in trajectory in the mean number of references per author after 2015.
4.3.2. Literature reviews are becoming more common
Literature reviews aim to capture the current state of a research area or field. Given that the production of scholarly articles is growing, we expect CHI authors are increasingly conducting literature reviews to address this expansion. We identified literature reviews by searching each CHI article’s full text (excluding the references section) for the keywords “literature review,” “literature survey,” and “systematic literature review.” We manually verified these occurrences, ensuring that only articles explicitly stating they conducted a literature review were included. Note that CHI authors, particularly in earlier proceedings, often refer to the related work section as a literature review. Additionally, some authors claim to have conducted an “extensive” or “exhaustive” literature review without detailing the process. Our data encompasses “literature review” in a broad sense, while “systematic literature reviews” are represented more narrowly, typically indicating articles that adhere to a more rigorous academic process.
@@ 4.3.2. Literature reviews are becoming more common
Figure 8. Absolute number (top) and relative percentage (bottom) of articles in the CHI proceedings reporting having conducted a “literature review” (including “literature surveys” and “systematic literature reviews”; in red) and a “systematic literature review” (in teal color) from 1981 to 2024. Systematic literature reviews in CHI articles are a new phenomenon starting in 2016.
[…]
A line plot of literature reviews in CHI proceedings from 1981 to 2024.
@@
In total, we identified 615 literature reviews and 60 systematic literature reviews. The time series in Figure 8 demonstrates that it is increasingly more common for CHI authors to conduct literature reviews. The amount of articles which mention having conducted a literature review was about 10% in the years 2021, 2022, and 2023. Figure 8 also demonstrates that systematic literature reviews in CHI are a recent phenomenon. Before 2016, there was no instance of authors calling their work a systematic literature review, but it has become increasingly more common for CHI authors to conduct systematic literature reviews. In 2024, the CHI proceedings included 17 systematic reviews (1.6% of the articles published in that year).

### v7

@@ 4. Findings
Over the course of 43 years (from 1981 to 2024, with exception of 1984), authors at the CHI Conference cited a total of 558,080 articles. The mean number of references per CHI article over the entirety of the CHI Proceedings is 48.35 references ($SD=35.09$, $Min=0$, $Max=496$). However, the field of HCI has evolved from its humble beginnings, and a more in-depth look at these descriptive statistics is warranted. In particular, we examine the increase in the mean number of references per CHI article in the following section.
4.1. The Rising Number of References in CHI Articles (RQ1)
We first provide a historical overview of citation trends in the CHI Proceedings, from the conference’s beginnings to the current trend. We then extrapolate this trend into the future to provide an outlook for the year 2030. This year is not too distant (only five years from the time of writing), and we believe focusing on the year 2030 in our prediction of future trends will – if the current trend is not broken – provide an accurate and realistic estimation of the future of citations at CHI.
@@ 4.1.1. Historical development
Figure 3. Excerpts from the Call for Papers (CfPs) at the CHI Conference in 2015 and 2016 (own highlighting).
In the early years of the CHI Conference, it was not uncommon for authors to submit an article without a single citation to other works. Zero-reference articles continued to be a part of the CHI Proceedings with only few exceptions until the year 1997 (see Figure 1 and Figure 2). From 1998 to 2017, the minimum number of references per article has been positive and in the one-digit range. Since 2017, the minimum number of references has increased into the double-digits.
Between 1981 and 2015, the mean number of references per CHI article showed only modest linear growth (cf. Figure 1 and Figure 2). After 2015, a clear and sudden quantitative change in the CHI community’s citation practices took place. Clearly, the CHI community entered a novel trajectory at this point in time, characterized by a year-by-year increase in the mean number of references per article. [^2] Throughout this article, we refer to the mean number of references. As depicted in Figure 1, the median number of references follows a very similar trend. While there were on average 28.8 references per article in the 2015 CHI Proceedings, the number of references per CHI article more than tripled to an average of 88.6 references per CHI article in 2024.
For about two decades since its inception, the maximum number of references per CHI article hovered at around 40. After 2015, the maximum number of references skyrocketed to up to 496 references in one article in 2022. The box plots in Figure 2 demonstrate that the maximum number of references per CHI article has become more volatile in recent years. During the years of the COVID-19 pandemic (2020 and 2021), one can observe a slump in the maximum number and upper quartile of the number of references per CHI article. However, there is, on average, a clear tendency for CHI authors to include more references in their articles compared to the years before 2015 and the mean number of references included in CHI articles is monotonously growing each year.
4.1.2. Extrapolation of the observed trend
We extrapolate the observed trend using simple linear regression models (LM1 and LM2; see Table 1) to predict the mean and minimum number of references per article in future instances of the CHI Conference Proceedings (see Figure 4). Both models are significant ($F(1,7)=308.96$, $p<1e^{-6}$ and $F(1,7)=6.81$, $p=0.03$, respectively), with model LM1 explaining 98% of the variance in the mean number of references. As depicted in Figure 4, the year-by-year growth in the mean number of references follows a near-linear pattern. Each year since 2016, the predicted mean number of references in CHI articles increases by approximately 6.45 references. We estimate that if the current observed trend is not broken, articles published at CHI 2030 will contain on average 129.5 references. The minimum number of references per article increases by 1.8 references per year. It is predicted to reach 33.7 references in 2030, which is more than the mean number of references from the year 2015 (28.9 references).
Table 1. Simple linear regression models evaluate the extent to which Year ($t$) can predict the mean (LM1) and minimum (LM2) number of References ($y$) in CHI articles. The two models are used to predict the number of references per article in future CHI proceedings in Figure 4.
LM1
(Mean number of references)
LM2
(Min number of references)
Year range 2016–2024 2016–2024
Regression equation $y={\beta_{0}}+{\beta_{1}}{\mathrm{t}}$ $y={\beta_{0}}+{\beta_{1}}{\mathrm{t}}$
Coefficient $\beta_{0}$ (intercept) -12971.92 -3620.33
Coefficient $\beta_{1}$ (slope) 6.45 1.80
t (for slope $\beta_{1}$) $t=17.58$
$p<1e^{-6}$
$t=2.61$
$p=0.03$
MAE 1.91 3.99
MSE 6.29 22.18
R2 0.98 0.49
F-test $F(1,7)=308.96$
$p<1e^{-6}$
$F(1,7)=6.81$
$p=0.03$
Figure 4. Extrapolation of the observed trend in the mean (red) and minimum (blue) number of references per CHI article. CHI articles are predicted to include on average almost 130 references in the year 2030. The minimum number of references per CHI article is estimated to reach 33.7 references per article in 2030, which is more than the 2015 average level of references.
4.1.3. Root cause
A diverse and complex community consisting of thousands of members, such as the CHI community, is not likely to suddenly change its course, unless change is mandated from top down. In our case, an editorial policy change launched the CHI community on its expansive path. In this section, we discuss the policy change that is the root cause for the observed non-linearity in the CHI community’s trajectory.
The CHI Conference organizing committee (Program Chairs) in consultation with the Executive Committee of the ACM Special Interest Group on Computer-Human Interaction (SIGCHI) introduced a decisive change in the year 2016. In this year, the conference’s Call for Papers (CfP) moved away from counting references toward the overall page limit of CHI articles, as highlighted in the excerpts reproduced in Figure 3. Articles published at CHI ’16 were allowed to include an unlimited number of references, as long as the article’s content fit the 10-page limit, with references no longer counting toward this page limit. While sub-committees were introduced as another important change in the year 2016, allowing an unlimited number of references clearly is the main cause of the increase in the mean number of references identified in our work.
Further, the “Guide to a Successful Paper or Note Submission” at CHI 2016 (ACM SIGCHI, 2016) stressed the importance of citing relevant previous work, explicitly asking “is prior work adequately reviewed?” The 2016 Guide also highlighted the importance of reproducibility of authors work, with “letting others build on your work” being regarded as “the entire purpose of a CHI Paper” (ACM SIGCHI, 2016). These direct quotes from the 2016 Guide were stark reminders of the importance of citing prior work and may have contributed to making the community pay special attention to their references in the year 2016.
These policy changes nudged the community on its expansive path. Several other factors, such as systematic bias in awards presented to articles, may have also played a role, as described in the following section.
4.2. Additional Contextual Factors Influencing HCI Citation Practices (RQ2)
The CHI community is vast and highly diverse, and there are many factors that potentially shape and make up the citation practices of the members of the CHI community. While the editorial policy change introduced in 2016 clearly is the enabler of the rise in the number of references included in CHI articles, in this section we take a step back and explore a number of contextual factors affecting the citation practices at the CHI Conference. The six factors we investigate are a potential bias in awards (Section 4.2.1), an increase in the average number of authors (Section 4.2.2), an increase in the number of literature reviews (Section 4.2.3), an increasing number of citations to arXiv pre-print articles (Section 4.2.4), an increase in citations to data and code repositories (Section 4.2.5), and an increase in the number of citations to potential predatory journals and publishers (Section 4.2.6). These six factors are used as confounding variables in our event study specification (see Section 4.3). We do not claim that this list of factors is exhaustive. Beside the investigated changes, other observable or latent factors could contribute to the trend, as is the case with any complex system under change.
4.2.1. Are high-citation articles rewarded in the CHI community?
Each year, awards are given to a select number of accepted articles at the CHI Conference. Typically, between 14.6% and 24.6% of articles are given an award at the CHI Conference (based on awards presented at CHI from 2015–2024). These awards are a signal to the community on which articles are to be considered high quality. As a quality signal, awards can motivate authors to change their behavior. If articles with a high number of references were to systematically be presented with awards, this could potentially incentivize CHI authors to include more references in their articles.
Table 2. Comparison of mean number of references in articles awarded a best paper award or honorable mention (Mawarded) with the mean number of references in unawarded articles (Mregular) in the CHI Proceedings from 2007 to 2024.
@@ 4.2.1. Are high-citation articles rewarded in the CHI community?
2007 29.966 23.368 2.860 [2.045, 11.149] 0.580 0.005 *
2008 28.029 22.598 2.632 [1.364, 9.499] 0.491 0.009 *
2009 28.106 23.739 2.453 [0.863, 7.872] 0.393 0.015 *
2010 29.138 24.181 2.999 [1.704, 8.21] 0.420 0.003 *
2011 25.972 26.582 -0.369 [-3.859, 2.64] -0.048 0.713
2012 27.923 27.069 0.496 [-2.534, 4.242] 0.068 0.620
2013 30.213 28.772 0.978 [-1.457, 4.339] 0.118 0.329
2014 31.398 28.66 2.036 [0.095, 5.381] 0.227 0.042 *
2015 31.219 28.153 2.298 [0.445, 5.685] 0.243 0.022 *
Policy Change
2016 43.470 40.705 1.616 [-0.597, 6.127] 0.170 0.107
2017 49.884 44.497 2.779 [1.581, 9.194] 0.283 0.006 *
2018 57.865 49.880 3.629 [3.665, 12.307] 0.369 $0$
2019 61.142 54.991 2.616 [1.534, 10.768] 0.242 0.009 *
2020 66.051 60.125 2.678 [1.582, 10.271] 0.241 0.008 *
2021 72.965 70.987 0.691 [-3.642, 7.598] 0.064 0.490
2022 87.922 81.933 1.520 [-1.747, 13.725] 0.150 0.129
2023 88.969 84.774 1.222 [-2.544, 10.934] 0.117 0.222
2024 92.958 87.651 1.795 [-0.494, 11.109] 0.144 0.073
Figure 5. The number of co-authors in CHI papers has increased over time, highlighting the trend toward increased collaboration in HCI.
Figure 6. Mean number of references per author in CHI articles
Figure 7. Mean number of references (primary axis) and mean number of authors (secondary axis) per CHI article.
@@
Table 2 shows no evidence of a significant systematic bias toward awarding high-reference articles in recent years (2021–2024). For instance, in the 2024 CHI Proceedings, the mean number of references for awarded articles was 92.96 compared to 87.65 for articles without award. This difference is not significant ($p>0.05$). However, awarded articles had significantly more references than unawarded articles ($p<0.05$; Cohen’s $d=0.23\ldots 0.37$) in the years immediately before and after the change (2014–2020, with exception of 2016). While awards were also given to high-reference articles in some earlier periods (2007–2010), awards were systematically given to high-reference articles around the policy change (with exception of 2016). This systematic awarding of articles with above-average references may have sent a quality signal to authors, thus potentially contributing to launching the CHI community on its current expansive trajectory.
@@ 4.2.2. Collaborative research is increasing
There has been an increase in collaboration in the field of HCI leading to articles with a greater number of co-authors. This is also reflected in the CHI Proceedings. Figure 5 depicts the number of authors per CHI article for each CHI proceedings year. The mean number of authors per CHI article has increased over time, from an average of 1.85 authors per article in 1981 to an average of 5.17 authors per article in 2024. Figure 5 also demonstrates that in the years between 2020 and 2024, there have been articles surpassing 20 authors in the CHI proceedings. The CHI 2024 Proceedings contain an article with 36 authors.
@@
Figure 7 plots the average number of references per author in CHI articles since the inception of the CHI Conference. In 1981, a CHI article included on average 5.6 references per author. This number stayed stable over three decades, with only a slight increase up to 2015. But there is a noticeable change in the mean number of references per author after 2015. Since 2016, spurred by the increase in the mean number of references per CHI article, the mean number of references per author has increased. In 2024, the mean number of references per author has reached 21.4 references per author.
Figure 7 plots the mean number of authors (right axis) against the mean number of references (left axis) per CHI article. While there clearly is a non-linearity visible in the latter time series, the former time series has been growing fairly linearly. The plot visually demonstrates that the growth in co-authorship cannot fully explain the growth in the mean number of references per CHI article.
@@
Figure 8. Absolute number of articles in the CHI proceedings reporting having conducted a literature review
Figure 9. Relative number of articles in the CHI proceedings reporting having conducted a literature review
Figure 10. Formal citations to arXiv pre-print papers in CHI articles
Figure 11. Relative number of articles formally citing arXiv pre-prints $n$ times
Figure 12. Mean number of citations to data and code repositories per CHI article
Figure 13. Relative number of CHI articles citing at least one data or code repositories
4.2.3. Literature reviews are becoming more common at CHI
In total, we identified 615 papers reporting having conducted a “literature review” and 60 systematic literature reviews in the CHI Proceedings from 1981–2024. The time series in Figure 9 showcases that since lifting the page restrictions in 2015, it has become increasingly common for CHI authors to conduct and publish literature reviews. About 10% of all articles in the latest four CHI Proceedings mention having conducted a “literature review,” with only a slight drop in 2024 (see Figure 9).
@@ 4.2.3. Literature reviews are becoming more common at CHI
Figures 9 and 9 demonstrate that systematic literature reviews in CHI are a recent and growing phenomenon. Before 2016, there was no instance of authors calling their work a systematic literature review. In recent years, it has become increasingly common for CHI authors to conduct systematic literature reviews, with up to 2% of the published articles in a given proceedings series reporting systematic literature reviews. In 2024, the CHI Proceedings included 17 systematic reviews (1.6% of the articles published in that year).
(a) (a) Absolute number of citations to potential predatory journals and publishers
(b) (c) Absolute number of citations to potential predatory journals and publishers (excluding Frontiers journals)
(c) (e) Absolute number of citations to Frontiers journals from CHI articles
(d) (b) Relative number of citations to potential predatory journals and publishers
(e) (d) Relative number of citations to potential predatory journals and publishers (excluding Frontiers journals)
(f) (f) Relative number of citations to Frontiers journals from CHI articles
Figure 14. Citations to potential predatory journals and publishers from CHI articles have become more common since 2016 (a and b). This growth is to a great extent produced by citations to large questionable publishers, such as Frontiers Media (e and f). While citations from CHI articles to smaller questionable publishers are growing in absolute terms (c), they remain limited in the recent decade in relative terms (d).